# DynamicKV: Task-Aware Adaptive KV Cache Compression for Long Context LLMs

### v2

@@ Abstract
Efficient KV cache management in LLMs is crucial for long-context tasks like RAG and summarization. Existing KV cache compression methods enforce a fixed pattern, neglecting task-specific characteristics and reducing the retention of essential information. However, we observe distinct activation patterns across layers in various tasks, highlighting the need for adaptive strategies tailored to each task’s unique demands. Based on this insight, we propose DynamicKV, a method that dynamically optimizes token retention by adjusting the number of tokens retained at each layer to adapt to the specific task. DynamicKV establishes global and per-layer maximum KV cache budgets, temporarily retaining the maximum budget for the current layer, and periodically updating the KV cache sizes of all preceding layers during inference. Our method retains only $1.7\%$ of the KV cache size while achieving $\sim 85\%$ of the Full KV cache performance on LongBench. Notably, even under extreme compression ($0.9\%$), DynamicKV surpasses state-of-the-art (SOTA) methods by 11% in the Needle-in-a-Haystack test using Mistral-7B-Instruct-v0.2. The code will be released.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Radford, 2018; Radford et al., 2019) are exerting a considerable influence in the field of natural language processing (NLP), driving advancements in summarization, translation, code generation, etc. (Chiang et al., 2023; Zhong et al., 2023; Peng et al., 2023; Lu et al., 2024; Wang et al., 2024). Recent developments in LLMs (Liu et al., 2024b) have been scaled up to handle long contexts, with LlaMA3 (Dubey et al., 2024) processing up to 32K tokens and InternLM (Cai et al., 2024) handling 1M tokens. However, scaling LLMs to handle extended contexts inherently incurs a substantial delay due to the quadratic complexity of attention mechanisms with increasing context length. A widely adopted solution to alleviate these delays is caching the key and value (KV) states of previous tokens (Waddington et al., 2013). Despite this optimization, handling long sequences still demands substantial memory (e.g., maintaining a KV cache for 100K tokens in LlaMA2-7B (Touvron et al., 2023) consumes over 50 GB of memory).
@@
Building on this insight, we propose a task-aware adaptive KV cache compression method, named DynamicKV. Specifically, we first calculate an attention score for the most recent few tokens and all other tokens, which in RAG (Lewis et al., 2020) can be viewed as calculating the relevance of the most recent query to the retrieved text. Then, we preset a temporary storage to hold the temporary KV cache states and gradually calculate the size of the final retained temporary storage at each k layer by calculating the size of the correlation mean. It should be noted that at each update, the value is gradually normalized, and the retained temporary storage at each layer is always smaller than the previous one. This temporary storage is determined by the number of tokens that need to be retained, and its size is much smaller than the original cache, thus imposing minimal memory overhead.
We validate our DynamicKV on 16 datasets from LongBench (Bai et al., 2023), demonstrating robust performance across multiple models, including LlaMA-3-8B-Instruct (Dubey et al., 2024), Qwen-2-7B-Instruct (Yang et al., 2024), Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), InternLM-2.5-7b-Chat-1M (Cai et al., 2024). Our DynamicKV exhibits superior overall effectiveness compared to conventional fixed-pattern methods (Zhang et al., 2024b; Li et al., 2024; Nawrot et al., 2024). Notably, DynamicKV can retain full performance while utilizing only 6.9% of the tokens, and in extreme scenarios, it preserves 90% of the performance with just 1.7% of the tokens. Furthermore, experiments on the Needle in a Haystack benchmark revealed that DynamicKV significantly outperforms state-of-the-art (SOTA) methods.
@@ 2 Related Work
Potential patterns of attention in LLMs. The Transformer architecture (Vaswani, 2017) becomes a cornerstone in NLP by stacking multiple layers to progressively refine input data. BERT (Devlin, 2018), a model based on this architecture, Jawahar et al. (2019) demonstrates that intermediate layers encode a rich hierarchy of linguistic information: from surface-level features at the bottom, through syntactic features in the middle, to semantic features at the top. This indicates that models are capable not only of understanding lexical information but also of grasping more complex linguistic structures.
For decoder-only LLMs, Fan et al. (2024) observes that not all layers are necessary for simple tasks, as intermediate layers can often achieve comparable performance to the final layer. Techniques like (Elhoushi et al., 2024), which involve increasing dropout in lower layers during training, allow the model to exit computation early, reducing resource consumption.
Figure 2: Average token retention across layers in LlaMA for different tasks, including (a) Single/Multi-Document QA, (b) Summarization, (c) Synthetic Task, and (d) Code Completion. There is a sharp decrease in token retention after the first layer, followed by varying patterns of fluctuation. Peaks are observed around Layer 15 and towards the final layers.
To optimize model inference efficiency, especially in terms of KV cache compression, Brandon et al. (2024) proposes cross-layer attention(CLA), which can reduce the KV cache size by at least half by sharing cross-layer attention, significantly lowering memory usage. Ada-KV (Feng et al., 2024) visualizes attention distributions across all layers have also shown that attention patterns dynamically evolve as the layers progress. Inspired by these findings, we aim to dynamically select and adjust the number of tokens to retain per layer, combining inter-layer redundancy identification with efficient KV cache management. This approach aims to maintain high-quality output while improving inference efficiency.
@@
Token drop is a strategy designed to reduce memory usage by selectively retaining the most influential tokens in the KV cache during the inference phase of LLMs. Due to its plug-and-play nature, the token drop method can often be applied to different models without incurring any additional costs. FastGen (Ge et al., 2023) evicts unnecessary contexts and discards non-special tokens based on the recognized structure of attention modules by effectively analyzing the token information within attention patterns. Scissorhands (Liu et al., 2024c) exploits the hypothesis of the persistence of importance, suggesting that tokens with significant influence at one point will continue to impact future generations. By using attention scores as a metric and applying a Least Recently Used (LRU) cache eviction strategy, it discards non-critical tokens to optimize memory usage. StreamingLLM (Xiao et al., 2023) leverages the characteristics of attention sinks in LLMs to focus on streaming processing with dynamic adjustment of the KV cache. H2O (Zhang et al., 2024b) proposes a scoring function based on accumulated attention scores for greedily evicting KV pairs during generation. SnapKV (Li et al., 2024) primarily achieves compression by selectively targeting key positions for each attention head. PyramidKV (Zhang et al., 2024a) identified the phenomenon of massive activation and adopted a hierarchical structure to optimize the number of KV cache entries retained at each layer. Although the PyramidKV approach considers the varying information density across different layers, its pyramidal pattern does not generalize across multiple models or tasks. LazyLLM (Fu et al., 2024) utilizes dynamic token pruning and an Aux Cache mechanism, allowing the model to select different subsets of tokens from the context at various generation steps, even reviving tokens pruned in previous steps. Ada-KV (Feng et al., 2024) breaks from the conventional approach of uniform budget allocation across attention heads within layers, and optimizes the eviction loss upper bound, leading to improved performance under various memory constraints when integrated with SnapKV and PyramidKV.
3 Observation
@@
To systematically investigate the attention mechanism across layers in LLMs for long-context inputs, we conduct a fine-grained analysis of four tasks: single- and multi-document question answering (QA), summarization, synthetic tasks, and code completion. The main target is to investigate the distribution of attention in these various tasks, thereby enhancing our understanding of how the model aggregates dispread information within long-context inputs to generate accurate responses.
@@
In particular, we focus our analysis on LlaMA (Dubey et al., 2024), visualizing the distribution and behavior of attention across layers to gain deeper insights into its internal mechanisms. Inspired by Zhang et al. (2024a), we calculate the average attention scores between the most recent tokens and all other tokens. Based on these scores, we then identify the top-k (128 multiplied by the number of layers) tokens with the highest attention across all layers, resulting in a layer distribution map denoted as Figure 2.
We observe a significant drop in the KV cache size requirement at the lower layers across the four tasks, indicating that only a small KV cache is needed in these layers. In contrast, the upper layers show a clear upward trend, suggesting that larger KV cache sizes are necessary, particularly in the code completion task, where complex reasoning is required. This phenomenon underscores that tasks involving complex reasoning demand larger KV cache sizes in the upper layers.
@@ 4 DynamicKV
During inference, the quadratic complexity of attention calculation results in a significant computational and memory burden, especially when processing long contexts. DynamicKV addresses this issue by focusing on inter-layer attention in large language models (LLMs), determining the appropriate size of KV cache to retain per layer through efficient awareness of inter-layer attention.
Rather than relying on a fixed retention pattern, such as pyramid-shape or average retention all layers, DynamicKV employs a progressive algorithm that dynamically adjusts token retention during the prefill phase. This dynamic retention strategy accelerates the decoding stage while maintaining minimal impact on overall memory usage.
Specifically, we first define layer $l\in\displaystyle\mathbb{R}^{L}$ and head $h\in\displaystyle\mathbb{R}^{H}$ in LLMs. For the calculation of attention scores, we use weights $W_{Q}\in\displaystyle\mathbb{R}^{N\times N}$, $W_{K}\in\displaystyle\mathbb{R}^{N\times N}$, and $W_{V}\in\displaystyle\mathbb{R}^{N\times N}$, with the input query embedding denoted as $X\in\displaystyle\mathbb{R}^{N\times M}$, $N$ is the dimension of the hidden size, and $M$ is the length of input tokens. Traditional token drop methods often consider the most recent tokens as the important ones for producing output information, as they retain relevant information needed for generating answers. We refer to these tokens collectively as the current window, with the window size denoted as $ws$. In the prefill phase, we adopt the method from Li et al. (2024), Zhang et al. (2024a), where the attention score is calculated by averaging over the current window and previous tokens, followed by pooling. The formula is as follows:
$A_{l,h}=\text{pooling}(\frac{1}{ws}\sum_{i=1}^{ws}\text{Attention}(X_{i},W_{Q},W_{K})),$ (1)
where pooling helps better understand the context and $A_{l,h}$ denotes the attention score for the l-th layer and the h-th head. This approach allows us to effectively pool the attention scores, ensuring that key tokens are retained based on their relevance to both the current window and previous context.
@@
Next, we set a fixed retention budget. Specifically, to ensure a fair comparison with other methods, we introduce the average retention length per layer, denoted as $wt$, and a scaling ratio, $r_{max}$. The calculation formula is as follows:
$bs=(wt-ws)\times r_{max},$ (2)
where $bs$ represents the size of retained KV cache across all layers. Next, we design a layer-aware progressive dynamic KV cache compression method. The prefill phase of LLMs involves a hierarchical forward process, where for each layer, we retain a KV cache of length $bs$ when computing $A$. Additionally, every $m$ layers, we perform an update across the current and all previous layers. Specifically, for each layer, we use a top-$k$ strategy to retain the largest $bs$ values from $A_{l}$, where $A_{l}$ represents the attention scores of layer $l$. The formula for this process is as follows:
$A_{l}^{\prime}=\text{TopK}(A_{l},bs).$ (3)
Next, we extract the indices in the original $A_{l}$ that correspond to the values in $A_{l}^{\prime}$. The KV cache at these indices is retained as the compressed KV cache. Specifically, the retained KV cache is defined as:
$\text{KV}_{l}^{\prime}=\text{KV}_{l}[A_{l}^{\prime}.\text{indices}],$ (4)
where $A_{l}^{\prime}.\text{indices}$ represents the indices of the top-$k$ values in $A_{l}$. This ensures that the KV cache is compressed efficiently, retaining only the most important tokens for each layer while minimizing memory usage.
To ensure that the memory required for hierarchical transmission remains small, the KV cache of each layer is initially compressed as described above. For every $m$ layer, we extract $A$ and perform a unified normalization across the completed layers, updating them layer by layer to ensure consistency across the entire hierarchy.
First, we fix the final size of the KV cache to be retained, which is calculated as $(wt-ws)\times H\times l$, where $H$ is the number of heads and $l$ is the number of layers. Then, for each layer, the attention score $A$ is used to compute the length to retain for each layer $C_{l}$ via a top-$k$ strategy. The retention lengths for the first $m$ layers are then normalized to obtain a budget length $Z$, ensuring that the retention is distributed effectively across layers. The specific formula is as follows:
@@
$C_{l}=\text{Norm}(\text{Count\_Elements}\left(I\right))$ (6)
$Z=\left[\frac{bs\times t}{\textit{max}(C_{l})}\ ~{}\textit{for}\ t\ \in\ C_{l}\right]$ (7)
$r=\frac{\sum Z}{(wt-ws)\times L},Z=\left[\frac{k}{r}\ ~{}\textit{for}\ k\ \in\ Z\right]$ (8)
The KV cache is further updated layer by layer based on this normalized budget, progressively refining the retained information to align with the overall compression strategy. The above process can be expressed as Algorithm 1.

### v3

@@ Abstract
Efficient KV cache management in LLMs is crucial for long-context tasks like RAG and summarization. Existing KV cache compression methods enforce a fixed pattern, neglecting task-specific characteristics and reducing the retention of essential information. However, we observe distinct activation patterns across layers in various tasks, highlighting the need for adaptive strategies tailored to each task’s unique demands. Based on this insight, we propose DynamicKV, a method that dynamically optimizes token retention by adjusting the number of tokens retained at each layer to adapt to the specific task. DynamicKV establishes global and per-layer maximum KV cache budgets, temporarily retaining the maximum budget for the current layer, and periodically updating the KV cache sizes of all preceding layers during inference. Our method retains only $1.7\%$ of the KV cache size while achieving $\sim 90\%$ of the Full KV cache performance on LongBench. Notably, even under extreme compression ($0.9\%$), DynamicKV surpasses state-of-the-art (SOTA) methods by 11% in the Needle-in-a-Haystack test using Mistral-7B-Instruct-v0.2. The code will be released.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023) are exerting a considerable influence in the field of natural language processing (NLP), driving advancements in summarization, translation, code generation, etc. (Chiang et al., 2023; Zhong et al., 2023; Peng et al., 2023; Lu et al., 2024; Wang et al., 2024). Recent developments in LLMs (Liu et al., 2024b) have been scaled up to handle long contexts, with LlaMA3 (Dubey et al., 2024) processing up to 32K tokens and InternLM (Cai et al., 2024) handling 1M tokens. Scaling LLMs to longer contexts introduces significant latency due to the quadratic complexity of attention. A common solution is to cache key and value (KV) status (Waddington et al., 2013), reducing computation. However, this comes at a high memory cost – for example, caching 100K tokens in LLaMA2-7B (Touvron et al., 2023) still requires over 50GB of memory.
@@
Building on this insight, we propose a task-aware adaptive KV cache compression method, named DynamicKV. Specifically, we first calculate an attention score for the most recent few tokens and all other tokens, which in RAG (Lewis et al., 2020) can be viewed as calculating the relevance of the most recent query to the retrieved text. Then, we preset a temporary storage to hold the temporary KV cache states and gradually calculate the size of the final retained temporary storage at each k layer by calculating the size of the correlation mean. It should be noted that at each update, the value is gradually normalized, and the retained temporary storage at each layer is always smaller than the previous one. This temporary storage is determined by the number of tokens that need to be retained, and its size is much smaller than the original cache, thus imposing minimal memory overhead. Experiments demonstrate that our DynamicKV can retain full performance while utilizing only 6.9% of the tokens, and in extreme scenarios, it preserves 90% of the performance with just 1.7% of the tokens. Furthermore, experiments on the Needle in a Haystack benchmark revealed that DynamicKV significantly outperforms state-of-the-art (SOTA) methods.
Contributions.
Our main contributions are:
  • We explore the impact of different task types on token retention at each layer of the LLM. Our findings highlight that for different tasks, token retention varies at each layer, and therefore, dynamic selection of token retention at each layer is necessary for different tasks.
  • Given our observation, we propose a novel KV cache compression method – DynamicKV to dynamically adjusts token retention during prefill phase.
  • Experimental results on the widely used long-context understanding benchmark, LongBench, demonstrate that our approach maintains full performance while using only 6.9% of the tokens.
@@ 2 Related Work
Potential patterns of attention in LLMs. The Transformer architecture (Vaswani, 2017) has driven progress in NLP through layered refinement of inputs. BERT (Devlin, 2018) reveals a hierarchical structure in intermediate layers via Jawahar et al. (2019): surface features dominate lower layers, evolving into syntactic and semantic representations toward the top. This underscores the capability of LLMs to encode both lexical and complex linguistic information across layers.
For decoder-only models, Fan et al. (2024) demonstrate that intermediate layers suffice for simple tasks, challenging the necessity of full-depth inference. Training strategies like (Elhoushi et al., 2024) further optimize efficiency by introducing layer-wise dropout, enabling early computation exit. Concurrently, KV cache optimization has emerged as a critical direction. Brandon et al. (2024) propose Cross-Layer Attention (CLA) to halve cache size via cross-layer attention sharing, while Feng et al. (2024) (Ada-KV) dynamically optimize eviction policies by analyzing cross-layer attention patterns. These works highlight the interplay between attention dynamics (Feng et al., 2024) and memory-efficient computation.
Figure 2: Analyzing the distribution of token retention across layers in LlaMA for different tasks, including Document QA, Summarization, Synthetic Task and Code Completion. (a) Each boxplot shows the distribution of token retention rates on different types of tasks across different layers. Results for different layers show that the token retention rates vary significantly across different tasks. (b) We visualize the token retention rates across different layers for four tasks, showing that the token retention rates exhibit different patterns across tasks.
@@
Token drop strategies for KV cache compression vary in approach but share a focus on identifying influential tokens. Attention-based methods like FastGen (Ge et al., 2023) and Scissorhands (Liu et al., 2024c) use attention patterns for pruning. Memory-aware approaches include StreamingLLM (Xiao et al., 2023), which prioritizes streaming via attention sinks, and H2O (Zhang et al., 2024b), which employs cumulative attention scoring for greedy eviction. Hierarchical methods like PyramidKV (Zhang et al., 2024a) adapt by layer but lack generalizability. SnapKV (Li et al., 2024) offers task-agnostic compression by selecting key positions per head. Dynamic frameworks such as LazyLLM (Fu et al., 2024) enable flexible token revival, and Ada-KV (Feng et al., 2024) improves overall performance by optimizing eviction loss bounds over uniform strategies.
Existing methods use fixed patterns across tasks, yet LLMs engage varying layers depending on the task Elhoushi et al. (2024). This suggests token retention during KV cache compression may also differ by task – an area largely unexplored. This paper examines how task type influences KV cache compression.
3 Preliminary Studies
@@
To systematically investigate the attention mechanism across layers in LLMs for long-context inputs, we conduct a fine-grained analysis on four different types of tasks: single- and multi-document question answering (QA), summarization, synthetic tasks, and code completion.
Experimental setting.
@@
In particular, we focus our analysis on LlaMA (Dubey et al., 2024), visualizing the distribution and behavior of attention across layers to gain deeper insights into its internal mechanisms. Inspired by Zhang et al. (2024a), we calculate the average attention scores between the most recent tokens and all other tokens. Based on these scores, we then identify the top-k (128 multiplied by the number of layers) tokens with the highest attention across all layers.
Observations.
As shown in Figure 2 (a), we use boxplot to visually present the distribution of four different types of tasks across different layers. We find that different tasks show significantly different token retention rates at a fixed layer. For example, at early layers, the spread is wide, indicating large task-specific variation. To further understand the distribution of token retention rates across different tasks, we visualize the token retention rates across all layers for each task, as shown in Figure 1 (b). We find that ❶ Synthetic Task shows higher retention rates in earlier layers, ❷ Code Completion shows higher retention rates in the earlier layers as well as the last three layers, and ❸ Document QA and Summarization exhibit different retention dynamics compared to others.
Insight.
The tokens to retain at each layer should adapt dynamically based on the task type.
@@ 4 DynamicKV
Previous work on KV cache compression Zhang et al. (2024a); Li et al. (2024) often allocaates a fixed KV cache size across LLM layers. However, as our analysis in § 3 demonstrates, attention patterns are not identical across different layers with different types of tasks. Therefore, using fixed KV cache size across layers on different tasks may lead to suboptimal performance. Thus, we propose DynamicKV— a dynamic layer-adaptive KV cache compression method. DynamicKV consists of two steps: (1) Dynamic Budget Allocation and (2) Progressive Cache Update.
4.1 Dynamic Budget Allocation
Traditional token drop methods often prioritize the most recent tokens, as these typically carry the most relevant context for generating the next output. We refer to this set of tokens as the current window, denoted by a window size $ws$. Tokens within this window are given the highest priority for retention. To manage memory efficiently, we first define a maximum KV cache retention budge per layer, denoted $B^{l}$, calculated as $B^{l}=(wt-ws)\times r_{max}$, where $r_{max}$ is a scaling ratio and $wt$ is the total number of tokens considered.
Following the approach of Li et al. (2024), we guide the selection of remaining tokens (outside the current window) based on their attention scores with respect to the instruction tokens. Tokens with higher attention scores are considered more relevant and are thus prioritized for retention in the GPU cache.
In a standard LLM, attention is computed as:
$A=softmax(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}),$ (1)
where $Q\in\mathbb{R}^{M\times d_{k}}$ and $K\in\mathbb{R}^{M\times d_{k}}$ are the query and key matrics, respectively, $d_{k}$ is the dimensionality of the key/queries, and $M$ is the sequence length. Inspired by Li et al. (2024); Zhang et al. (2024a), we compute per-layer attention scores $A^{l}$ over the current window using a multi-head pooling operation:
$A^{l}=Pooling(A[:,ws]).$ (2)
We then select the top $B^{l}$ tokens based on the highest values in $A^{l}$. The corresponding KV states at these positions are retained to form a compressed cache:
$KV_{retained}^{l}=KV^{l}[arg\ topK(A^{l},B^{l})].$ (3)
4.2 Progressive Cache Update
To further reduce KV cache usage in the middle layers, we partition the model into blocks of $m$ consecutive layers. For each such block, we dynamically determine the minimal initial retention threshold required to meet cumulative retention demands, while also refreshing the historical KV cache. At the end of each $m$-layer block, we normalize the retention scores to prioritize operationally critical tokens. This process yields a layer-specific budget allocation $Z^{\prime}$, which facilitates an efficient and adaptive distribution of the cache budget across layers. Specifically, we apply a top-K selection to retain the most relevant tokens across these layers, and the compute the retention count per layer using a counting function $\Phi$:
$C^{l}=Norm(\frac{1}{n}\cdot\Phi(TopK(A,(wt-ws)\times n),$ (4)
where $n$ is the number of progressive update layers processed so far, and $(wt-ws)$ denotes the number of tokens outside the current window.
Next, we compute a provisional budget $Z$ by scaling each layer’s retention score relative to the maximum:
$Z=\left[\frac{B^{l}\times t}{\textit{max}(C^{l})}|\ t\in C^{l}\right],$ (5)
where $B^{l}$ is the per-layer retention budget. This is then normalized across layers to ensure the total budget $B=(wt-ws)\times L$ is respected:
$Z^{\prime}=[k\cdot\frac{B}{\sum Z}|k\in Z].$ (6)
@@ 4.2 Progressive Cache Update
In practice, during the progressive update of the first $m$ layers, the mechanism uses the attention scores $A$ to estimate the optimal number of tokens to retain per layer. The function $\Phi$ counts the top-K attention entries assigned to each layer, forming $C^{l}$, which is then normalized into $Z$. Finally, the budget $Z^{\prime}$ governs how the KV cache is refined for each layer, enabling an adaptive and effective compression strategy across the different layers.
The above process can be expressed as Algorithm 1.
Algorithm 1 DynamicKV in Prefill Phase
1:  Input: initial budget K/V cache list ${K}^{b}$, ${V}^{b}$, ratio max $r_{max}$, update interval $m$, mean token length $wt$, window size $ws$, sequence length $S$, head dimention $d_{k}$, input embedding of window size ${X}^{ws}\in\mathbb{R}^{ws*d}$, initial budget Attention list computed by window token and others ${A}^{b}$,
2:  Output: Compressed K/V cache ${K}^{c}$, ${V}^{c}$
3:  $B^{l}=(wt-ws)\times r_{max}$
4:  def Update_Buffer_Length($A$, $l$):
5:       ${A}^{gather}$leftarrow cat(([${A}$ for $l$ in (1, $l$)]), 0).view(-1)
7:       Compute the $norm$ of $cnts$, range in (0, 1)
8:       $Z$leftarrow [int(($B^{l}*t$ / max($norm$))) for $t$ in $norm$]
9:       $r$leftarrow sum($Z$) / (($wt-ws$)$*L$)
10:       $Z^{\prime}$leftarrow [int($k/r$) for $k$ in $Z$]
11:       Return $Z^{\prime}$
12:  for $l\leftarrow 1\textbf{ to }L$ do
13:     Compute full KV states ${K}^{s}$, ${V}^{s}$
14:     for $h\leftarrow 1\textbf{ to }H$ do
16:        ${A}^{l}_{h}$leftarrow softmax((${X}^{ws}{W}_{h}^{Q}$) $\cdot$ ${K}_{h}^{T}$).mean(dim=-2).pooling(dim=-1)
17:     end for
18:     Append ${A}^{l}$ to ${A}^{b}$ /* current ${A}_{l}$ shape is [$H$, $S$] */
19:     /* calculate current layer buffer KV cache */
20:     indices $\leftarrow$ ${A}^{l}$.topk($B^{l}$, dim=-1).indices.unsqueeze(-1).expand(-1, -1, $d_{k}$)
23:     /* gradually compress*/
@@
25:        $Z^{\prime}$leftarrow Update_Buffer_Length(${A}^{l}$, $l$)
26:        /* update the buffer K/V Cache*/
27:        for $i$leftarrow 1 to $l$ do
28:           ${K}_{i}^{b}$leftarrow cat((${K}_{l}^{b}$[:,:${Z^{\prime}}_{i}$,:], ${K}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
29:           ${V}_{i}^{b}$leftarrow cat((${V}_{l}^{b}$[:,:${Z^{\prime}}_{i}$,:], ${V}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
30:        end for
31:     end if
32:  end for
33:  Update the K/V Cache ${K}^{c},{V}^{c}$ from ${K}^{b},{V}^{b}$
4.3 Implementation Details
Durint the inference, the process is divided into two phases, the prefilling phase and the decoding phase, consistent with existing inference engines Kwon et al. (2023). Our DynamicKV, while potentially encountering sample-specific attention patterns when determining the optimal KV cache size per layer, performs this step during the prefilling phase. During the decoding phase, no modifications are applied.
Q1: Does the DynamicKV handles batched inference?
A1: Yes. In fact, modern LLM inference and serving engines (e.g., vLLM Kwon et al. (2023)) generally process samples individually (i.e., batch size=$1$) in prefilling phase, while decoding allows for efficient parallel computation in batches. Since our DynamicKV introduces no modifications during decoding, our method aligns seamlessly with existing inference engines, ensuring that the decoding phase remains fully compatible with batched execution for high-throughput generation.
Q2: How does the DynamicKV compatible with FlashAttention?
A2: Our DynamicKV can compatible with FlashAttention during the decoding phase. Although our DynamicKV modifies the computation of attention scores during the prefilling phase, which limits compatibility with FlashAttention, it remains highly efficient. This is because attention is computed only within a small widow size $ws$, where $ws\ll M$, keeping the overhead minimal even without FlashAttention. In contrast, no modifications are applied in decoding phase, where we take advantage of FlashAttention to significantly improve computational efficiency.